OpenStudy (anonymous):

What is the equation of a line passing through (-3, 7) and having a slope of -1/5? A. y=-5x+22 B. y=5x+22 C. y=1/5x + 32/5 D. y=-1/5x+32/5 E. y=-1/5x+22

pooja195 (pooja195):

We use the point slope formula \[\huge~\rm~y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (triciaal):